# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

### v1

@@ Abstract
Dwarf scale cores— Ref. Zhang et al. (2025b) finds that diffuse dwarf galaxies exhibit unexpectedly stronger large-scale clustering than denser ones. This can be naturally explained in SIDM, where halos in the core-forming phase develop larger cores if they evolve longer. Since halos in overdense regions form earlier, they are more likely to host diffuse galaxies, leading to enhanced clustering. To demonstrate that our proposed two-component models can reproduce this behavior, we simulate isolated halos with mass $M=10^{11}~{}\rm M_{\odot}$, which host the dwarfs in Ref. Zhang et al. (2025b), using SIDM2c. These simulations show that the halos develop time-dependent cores, similar to those in the one-component model with $\sigma/m=0.3~{}\rm cm^{2}/g$ (SIDM03), as favored in that work.
Figure 2 compares the resulting density profiles at $t=5~{}$Gyr for halos in SIDM03 (dashed), and SIDM2c (solid), with median (green), $+1\sigma$ (orange), and $-1\sigma$ (blue) concentrations, based on the redshift zero $c_{200}-M$ relation from Ref. Dutton and Macciò (2014). We find that both the one- and two-component models produce similarly sized cores, confirming that SIDM2c, and by extension its velocity-dependent counterpart SIDM2v, retains the ability to explain dwarf galaxy clustering despite its richer internal structure.
@@
The inset panel in Fig. 2 tracks the time evolution of central densities. We observe that more concentrated halos evolve more rapidly in the two-component case, reaching the minimum central density at earlier times compared to their one-component counterparts. This suggests that mass segregation enhances the efficiency of gravothermal evolution in dense systems, driving the most extreme cases into core collapse within a Hubble time. This has important implications for the origin of the highly compact dark substructures inferred from strong lensing Vegetti and Koopmans (2009); Vegetti et al. (2010, 2012); Hezaveh et al. (2016); Minor et al. (2021); Tajalli et al. (2025). A small population of high-concentration dwarf halos can evolve into dense, collapsing configurations even with modest self-interaction cross sections that are insufficient to induce core collapse in one-component SIDM, as explored below.
@@
Figure 4: Comparison of simulated gravitational lenses in CDM (left), SIDM2v (middle), and SIDMx (right). Top panels show tangential (opaque) and radial (40% transparent) critical curves overlaid on projected density maps; bottom panels show corresponding caustics. The two-component models produce significantly more secondary caustics due to dense substructures. The GGSL cross section ($\sigma_{\rm GGSL}$) is computed summing over the areas enclosed by the secondary caustics whose effective Einstein radii are larger than $0.5^{\prime\prime}$ and residing within the $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo. The SIDM2v has $\sigma_{\rm GGSL}=6.76$, being about $13$ times that of the CDM case and aligns with the observation by Ref. Meneghetti et al. (2020). The SIDMx model represents an extreme case with only inter-species interaction and $\sigma_{\rm GGSL}=12.8$.
@@
Strong lensing perturber— SDSSJ0946+1006 provides a pronounced example of a compact, dark substructure revealed through strong lensing, with image reconstruction indicating a steep projected density slope, $\gamma_{\rm 2D}\sim-1.25$, and high mass within the inner kiloparsec Minor et al. (2021). Such subhalos challenge both CDM and single-component SIDM models. In CDM, it is extremely difficult to generate sufficiently dense and compact subhalos without invoking rare statistical fluctuations. While core-collapsing SIDM models can, in principle, produce such dense objects, prior studies have shown that cross sections of $\gtrsim 100~{}\rm cm^{2}/g$ are required for a $10^{11}\,M_{\odot}$ halo to reach collapse Nadler et al. (2023); Li et al. (2025), creating tension with the much smaller cross sections inferred from dwarf galaxy clustering. Moreover, halos that reach such densities would typically be expected to host observable galaxies, in conflict with the non-detection of luminous counterparts.
@@
Figure 3 shows the evolution of the benchmark halos in the plane of projected inner mass (within 1 kpc) versus average projected slope $\gamma_{\rm 2D}$, with arrows indicating their trajectories over time. All three evolve toward the region favored by the SDSSJ0946+1006 perturber, with M08 passing through the 95% confidence contour at $t\approx 7.4$ Gyr. This demonstrates that mass segregation in two-component SIDM can drive certain halos into dense, lensing-relevant configurations even at modest cross sections. By contrast, in SIDM03, the same halos remain in the core-forming phase and evolve away from the dense regime required to match the perturber. To connect with the cosmological context, we overlay subhalos from the SIDM2v simulation. Despite resolution limits, these subhalos extend to lower slopes and higher inner masses, with several candidates residing within the region favored by the SDSSJ0946+1006 perturber, supporting the viability of forming compact, dark structures in SIDM models with mass segregation.
@@
Small-Scale Lenses and Lensing Cross Sections— The capability of enhancing inner masses in two-component models also makes subhalos more efficient lenses and addresses the challenge reported in Ref. Meneghetti et al. (2020), where the GGSL cross section, defined as the area enclosed by tangential caustics induced by subhalos, is an order of magnitude larger in observations. This challenge remains difficult to reconcile with current models. Simulations with baryonic physics can increase subhalo densities, but the effect is often driven by one or two massive satellites and fails to robustly match observed statistics Bahé (2021); Robertson (2021); Dutra et al. (2025). In SIDM, the core collapse has demonstrated potential, but achieving it across all subhalos with inner density slopes steeper than $r^{-2.5}$ remains challenging Dutra et al. (2025). Moreover, when baryons are included, core-collapsed subhalos tend to enhance radial caustics rather than the tangential ones relevant for four-image systems Yang and Yu (2021). In contrast, mass segregation naturally increases the central densities and steepens the inner density slopes of subhalos. To facilitate comparison with the GGSL measurment in Ref. Meneghetti et al. (2020), we fix the redshift of our host system at $z=0.439$ and compute lensing properties by including all subhalos within a $5~{}{\rm Mpc}/h$ sphere. The lensing potential and deflection angles are primarily modeled analytically using the parametric SIDM halo profile developed in Hou et al. (2025); Yang et al. (2024), which allows for efficient calculation of lensing observables across a wide range of halo masses and internal structures. See the Supplemental Material for details.
[…]
@@
Conclusion— Our simulations show that SIDM with mass segregation retains cores in most dwarf halos, accelerates inner condensation in high-concentration halos, and yields denser subhalos in clusters. These results highlight SIDM2v as a unified and testable framework for resolving multiple small-scale challenges. Looking ahead, our framework may also shed light on the formation of early massive galaxies and supermassive black holes, including the so-called “little red dots”, as mass segregation remains effective at high redshifts. Our results suggest that small-scale structure may already be revealing not only the strength but also the multi-component nature of dark matter.
@@ Acknowledgements.
We thank Moritz Fischer, Guoliang Li, Nan Li, Ran Li, Shihong Liao, Priyamvada Natarajan, Huiyuan Wang, Hai-Bo Yu, and Yi-Ming Zhong for helpful discussion. The authors were supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), Chinese Academy of Sciences (CAS), and the Project for Young Scientists in Basic Research of CAS (No. YSBR-092).

### v2

@@ IV Dwarf scale cores
Ref. Zhang et al. (2025b) finds that diffuse dwarf galaxies exhibit unexpectedly stronger large-scale clustering than denser ones. This can be naturally explained in SIDM, where halos in the core-forming phase develop larger cores if they evolve longer. Since halos in overdense regions form earlier, they are more likely to host diffuse galaxies, leading to enhanced clustering. To demonstrate that our proposed two-component models can reproduce this behavior, we simulate isolated halos with mass $M=10^{11}~\rm M_{\odot}$, which host the dwarfs in Ref. Zhang et al. (2025b), using SIDM2c. These simulations show that the halos develop time-dependent cores, similar to those in the one-component model with $\sigma/m=0.3~\rm cm^{2}/g$ (SIDM03), as favored in that work.
We simulate the median-concentration halo with and without a collisionless stellar component, which follows median stellar-to-halo mass and size-mass relations. Figure 3 (left) presents results in the same style as Fig. 2 (right). While the halo densities in the baryonic cases (solid) are systematically higher than the DM-only cases (dashed), we found two crucial differences. One is that the stellar component develop cores in time, a feature directly addresses the observations. The other is that halos also have growing cores alongside with the increasing inner densities. This feature is precisely the opposite of the dark matter only case, where the core formation is accompanied with a decreasing inner halo density.
To track core growth clearly, we fit the halo density profiles with a $\beta 4$ model, defined as $\rho_{\beta 4}=\rho_{s}/((r/r_{s})^{4}+(r/r_{c})^{4})^{1/4}/(1+r/r_{s})^{2}$ Gilman et al. (2021); Yang et al. (2024), and plot the resulting core size $r_{c}$ evolution in Fig. 3 (right). After including baryons, the core size is significantly reduced due to gravitational contraction. Crucially, the core growth in SIDM2c appears close to that of the one-component SIDM03 after incorporating baryons.
In Fig. 4, we present the distribution of the gravothermal phase $\tau$ from parametric model predictions to show the dominance of cored halos in the dwarf halo population. The sample is selected by considering halos with masses higher than $10^{10}~\rm M_{\odot}/h$ and within 5 Mpc/h of the cluster host halo in the cosmological zoom-in simulation. For dwarf clustering analysis, we consider isolated halos (Group B, light blue) and therefore exclude cluster subhalos (i.e., those residing within the virial radius of the cluster halo). These halos are predominantly core forming, as their $\tau$ values are low, with mean and standard deviation both close to 0.2 Yang et al. (2024). The cluster subhalos (Group A, light red) have higher $\tau$ values because their progenitor halos are more massive. The mean of the distribution is shifted to 0.38, and there is a small population of halos that have undergone deep core collapse ($\tau>1$).
@@
Figure 6: Comparison of simulated gravitational lenses in CDM (left), Model-2v (middle), and Model-x (right). Top panels show tangential (opaque) and radial (40% transparent) critical curves overlaid on projected density maps; bottom panels show corresponding caustics. The two-component models produce significantly more secondary caustics due to dense substructures. The GGSL cross section ($\sigma_{\rm GGSL}$) is computed summing over the areas enclosed by the secondary caustics whose effective Einstein radii are larger than $0.5^{\prime\prime}$ and residing within the $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo.
V Strong lensing perturbers
We have seen that mass segregation can produce higher inner halo densities than one-component SIDM, even during core formation. A denser inner halo also accelerates the halo’s gravothermal evolution. This is intuitive from the dependence of the core-collapse time $t_{c}$ on halo concentration $c$. A dimensional analysis of $t_{c}$, translating the NFW parameters into mass and concentration, yields $t_{c}\propto c^{(a-7)/2}$ for a generic $v^{-a}$ cross section in one-component SIDM. For both a velocity-independent case ($a=0$) and the commonly used Coulomb-like scattering ($a=4$), $t_{c}$ decreases sharply as $c$ grows. Hence, a more compact halo is expected to have a shorter $t_{c}$.
Quantitatively, we calibrate the core collapse time of halos simulated in the SIDM2c model and find that it is close to that of a one-component SIDM model with $\sigma/m\approx 14~\text{cm}^{2}/\text{g}$, rather than the $0.3~\text{cm}^{2}/\text{g}$ value discussed earlier. As shown in Fig. 4, a small population of high concentration dwarf halos can evolve into dense, collapsing configurations under SIDM2c, explaining the observed dense dark substructures in strong lensing systems, commonly referred to as strong lensing perturbers Vegetti and Koopmans (2009); Vegetti et al. (2010, 2012); Hezaveh et al. (2016); Minor et al. (2021); Tajalli et al. (2025); Kong et al. (2025); Powell et al. (2025).
@@ V Strong lensing perturbers
Observationally, SDSSJ0946+1006 provides a pronounced example of a compact, dark substructure revealed through strong lensing, with image reconstruction indicating a steep projected density slope, $\gamma_{\rm 2D}\sim-1.25$, and high mass within the inner kiloparsec Minor et al. (2021). Such subhalos challenge both CDM and single-component SIDM models. In CDM, it is particularly challenging to generate sufficiently dense and compact subhalos without invoking rare statistical fluctuations. While core-collapsing SIDM models can, in principle, produce such dense objects, prior studies have shown that cross sections of $\gtrsim 100~\rm cm^{2}/g$ are required for a $10^{11}\,M_{\odot}$ halo to reach collapse Nadler et al. (2023); Li et al. (2025), creating a potential tension with the much smaller cross sections inferred from dwarf galaxy clustering. This tension could be alleviated by incorporating a $v^{-4}$ velocity-dependence into the SIDM model, which would allow lower-mass halos to have larger cross sections, enabling some $10^{9}~\mathrm{M}_{\odot}$ halos to undergo core collapse.
@@
Figure 5 shows the evolution of the benchmark halos in the plane of projected inner mass (within 1 kpc) versus average projected slope $\gamma_{\rm 2D}$, with arrows indicating their trajectories over time. All three evolve toward the region favored by the SDSSJ0946+1006 perturber, with M08 passing through the 95% confidence contour at $t\approx 7.4$ Gyr. This demonstrates that mass segregation in two-component SIDM can drive certain halos into dense, lensing-relevant configurations even at modest cross sections. By contrast, in SIDM03, the same halos remain in the core-forming phase and evolve away from the dense regime required to match the perturber. To connect with the cosmological context, we overlay subhalos of the cluster host in SIDM2c, obtained using the parametric model. These subhalos extend to lower slopes and higher inner masses, with several candidates residing within the region favored by the SDSSJ0946+1006 perturber, supporting the viability of forming compact, dark structures in SIDM models with mass segregation.
VI Small-Scale Lenses and Lensing Cross Sections
@@
The capability of enhancing inner masses in two-component models makes subhalos more efficient lenses and addresses the challenge reported in Ref. Meneghetti et al. (2020), where the GGSL cross section, defined as the area enclosed by tangential caustics induced by subhalos, was initially found to be an order of magnitude larger in observations. Refs. Meneghetti et al. (2022, 2023) further explored the effects of simulation resolution, AGN feedback, and galaxy formation models, etc., finding that the observed GGSL cross section remains a persistent factor of 3–6 times higher than predicted. This challenge remains difficult to reconcile with current models. Simulations with baryonic physics can increase subhalo densities, but the effect is often driven by one or two massive satellites and fails to robustly match observed statistics Bahé (2021); Robertson (2021); Dutra et al. (2025). In SIDM, the core collapse has demonstrated potential, but achieving it across all subhalos with inner density slopes steeper than $r^{-2.5}$ remains challenging Dutra et al. (2025). Moreover, when baryons are included, core-collapsed subhalos tend to enhance radial caustics rather than the tangential ones relevant for four-image systems Yang and Yu (2021). In contrast, mass segregation naturally increases central densities, raising the enclosed mass above the critical surface density and enlarging the Einstein radii of SIDM halos.
To quantify our framework’s capability to reconcile the GGSL anomaly, we perform cosmological simulations for CDM, SIDM2v, and SIDMx. The CDM simulation assumes a one-component model, while the latter two are detailed in Table 1. The resolution of these simulations is insufficient to make accurate predictions for dwarf halos. Therefore, we adopt a hybrid approach. We use simulated halo density profiles only for candidates with masses above $10^{13}~\rm M_{\odot}$ in the CDM simulation. These massive halos are matched across all three simulations, and only matched halos are retained, ensuring that the large-scale densities remain consistent with the respective cosmological runs. For lower-mass halos, we apply the parametric SIDM2c model to the corresponding CDM halos. We adopt the integral approach from Ref. Yang et al. (2024) to obtain predictions that incorporate the effects of accretion histories. This approach provides results that differ from the simulated SIDM2v and SIDMx models. For clarity, we refer to these hybrid models as Model-2v and Model-x, corresponding to cases where the massive halos are taken from the SIDM2v and SIDMx simulations, respectively. In effect, our study focuses on the impact of SIDM2c halos in the mass range $10^{10}$-$10^{12}~\rm M_{\odot}$ on GGSL. The integrated results are presented for two alternative models: one representing a scenario that remains safe from cluster SIDM constraints (Model-2v), and the other exhibiting extreme mass segregation (Model-x). We also incorporate the effect of baryons on halo density profiles using a parametric approach, obtaining results for the Model-2v+baryons and Model-x+baryons scenarios. For consistency with the CDM comparisons, the baryon potentials themselves are not included directly in the lensing analysis.
To facilitate comparison with the GGSL measurment in Ref. Meneghetti et al. (2020), we fix the redshift of our host system at $z=0.439$ and compute lensing properties by including all subhalos within a $5~{\rm Mpc}/h$ sphere. The lensing potential and deflection angles are primarily modeled analytically using the parametric SIDM halo profile developed in Hou et al. (2025); Yang et al. (2024), which allows for efficient calculation of lensing observables across a wide range of halo masses and internal structures. See the Supplemental Material for details.
Figure 6 compares the tangential (opaque) and radial (40% transparent) critical curves (top panels) and associated caustics (bottom panels) in the CDM, Model-2v, and Model-x simulations. We observe a substantial increase in the number and size of small-scale caustics in both two-component SIDM models compared to CDM.
Quantitatively, we compute the fiducial GGSL cross section ($\sigma_{\rm GGSL}$, in unit of arcsec2) as the total area enclosed by secondary caustics (excluding the host caustic), within a $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo. Following Meneghetti et al. Meneghetti et al. (2020), we apply a lower area threshold of $\pi(0.5^{\prime\prime})^{2}$. We also exclude any contribution with effective Einstein radius $\theta_{E}\equiv\sqrt{A_{c}/\pi}$ being larger than three arc-second, where $A_{c}$ is the area enclosed by a critical line. This cut prevents the statistical dominance of massive outliers, as they are physically distinct from dwarf halos and are not observed to produce $\theta_{E}>3^{\prime\prime}$ Meneghetti et al. (2022, 2023). We repeated the lensing analysis 11 times, corresponding to 11 different viewing angles. Each viewing angle was generated by consecutively rotating the system around the $x$- and $y$-axes in steps of $18^{\circ}$, i.e., $18^{\circ}$, $36^{\circ}$, $54^{\circ}$, $72^{\circ}$, and $90^{\circ}$ rotations for each axis, plus a case without rotation.
Table 2: GGSL cross section comparisons. The mean GGSL cross sections with their standard deviations and total number of contributing secondary caustics (aggregated over 11 analyses) are reported for different model scenarios. The ratio column computes the ratio of mean GGSL relative to the value in CDM, which illustrates the amount of excess. The GGSL cross section is computed considering secondary critical lines with $\theta_{E}<3^{\prime\prime}$.
Model Mean GGSL Std Dev Total N Ratio
CDM 0.3024 0.2484 41 1.00
Model-2v 0.6133 0.3689 85 2.03
Model-x 2.223 0.7615 207 7.35
Model-2v+baryons 1.360 0.5023 191 4.50
Model-x+baryons 4.149 1.602 326 13.7
Table 2 summarizes the results of our GGSL analysis. The table presents, for each simulation, the mean GGSL cross section, its standard deviation, and the total number of contributing secondary caustics (aggregated over 11 analyses). The final column reports the ratio of the mean GGSL cross section relative to the CDM value, quantifying the amount of excess produced by two-component SIDM models. All values are computed considering only secondary critical lines with $\theta_{E}<3^{\prime\prime}$. The Model-2v and Model-x benchmarks correspond to host halos that are shallower and denser, respectively, than the CDM host halo. As expected, the GGSL cross section is higher in the denser Model-x case. In Model-2v, the GGSL cross section and the number of contributing caustics are both approximately double the CDM values. This enhancement increases to about 4.5 times the CDM value when baryons are included. The excess is systematically larger in Model-x, where the ratio rises to 7.35 without baryons and 13.7 with baryons. The level of GGSL excess produced by our benchmark models demonstrates their compelling capability to explain the observed GGSL excess, which is a factor of 3-6 above theoretical expectations.
VII Conclusion
Observations on small scales reveal the simultaneous presence of cored dwarf galaxies and dense substructures in host halos spanning multiple mass scales. This study demonstrates that such features naturally coexist in SIDM models with mass segregation. We also reveal a conditioned universality in two-component SIDM models, which we use to construct a parametric model for generating theoretical predictions for a population of halos.
Future work could improve upon several aspects of this work, including the simulation resolution, the scan of model parameters, and the incorporation of baryonic physics. While controlled simulations are used to illustrate the effect of baryons, more dedicated simulations are required to map the full landscape of baryonic effects in SIDM2c halos. Moreover, the present study is limited to a single cluster system, which complicates direct comparison with observations.
@@ VII Conclusion
Despite these limitations, our current results highlight SIDM with mass segregation as a unified and testable framework for resolving multiple small-scale challenges. Looking ahead, our framework may also shed light on the formation of early massive galaxies and supermassive black holes, including the so-called “little red dots”, as mass segregation remains effective at high redshifts. Our results suggest that small-scale structure may already be revealing not only the strength but also the multi-component nature of dark matter.
@@ Acknowledgements.
We thank Moritz Fischer, Guoliang Li, Nan Li, Ran Li, Shihong Liao, Massimo Meneghetti, Priyamvada Natarajan, Huiyuan Wang, Hai-Bo Yu, and Yi-Ming Zhong for helpful discussion. The authors were supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), Natural Science Foundation of China under No. 12588101, Chinese Academy of Sciences (CAS), and the Project for Young Scientists in Basic Research of CAS (No. YSBR-092).